# Inferring Outcome Means of Exponential Family Distributions Estimated by Deep Neural Networks

### v2

@@ 4 Numerical Experiments
$n=400,r=n^{0.8}$0.69 0.75(0.59) -0.19 0.21(0.33) 0.95 0.70 85.6% 0.63(0.60)
$r=n^{0.85}$0.46 0.59(0.51) -0.15 0.19(0.31) 0.78 0.70 93.3% 0.78(0.75)
$r=n^{0.9}$0.27 0.51(0.44) -0.11 0.18(0.30) 0.67 0.69 95.6% 0.94(0.93)
Boostrap with $f_{0}$0.19 0.72(0.59) -0.07 0.25(0.36) 0.93 - 94.3% 2.25(2.39)
$n=700,r=n^{0.8}$0.40 0.59(0.50) -0.14 0.19(0.31) 0.78 0.67 93.8% 0.75(0.68)
$r=n^{0.85}$0.17 0.50(0.42) -0.09 0.18(0.29) 0.66 0.67 97.0% 0.93(0.88)
$r=n^{0.9}$ 0.01 0.47(0.39) -0.04 0.17(0.27) 0.61 0.65 95.9% 1.07(1.05)
Boostrap with $f_{0}$ 0.04 0.70(0.56) -0.01 0.25(0.34) 0.90 - 96.1% 2.49(2.61)
Binomial Model
$n=400,r=n^{0.8}$0.21 0.58(0.54) -0.03 0.09(0.10) 0.79 0.60 89.7% 0.33(0.01)
$r=n^{0.85}$0.14 0.57(0.51) -0.02 0.09(0.10) 0.76 0.62 91.1% 0.35(0.02)
$r=n^{0.9}$0.08 0.56(0.49) -0.02 0.09(0.10) 0.75 0.63 91.1% 0.36(0.02)
Boostrap with $f_{0}$0.28 0.90(0.76) -0.02 0.13(0.13) 1.18 - 96.9% 0.69(0.22)
$n=700,r=n^{0.8}$0.18 0.47(0.45) -0.02 0.07(0.09) 0.66 0.54 91.1% 0.29(0.01)
$r=n^{0.85}$0.12 0.45(0.43) -0.02 0.07(0.08) 0.62 0.55 92.9% 0.30(0.02)
$r=n^{0.9}$0.06 0.44(0.42) -0.01 0.07(0.08) 0.60 0.56 94.2% 0.32(0.02)
Boostrap with $f_{0}$0.03 0.60(0.50) -0.00 0.10(0.10) 0.78 - 95.0% 0.44(0.16)
Figure 2: Histograms of $\psi^{\prime}(\widehat{f}^{B}(x_{\text{test}}))$ and $\psi^{\prime}(f_{0}(x_{\text{test}}))$. Results are shown for 2(a) logistic model, 2(b) Poisson model (count outcomes) and 2(c) Binomial model (count outcomes) with $n=700$ and subsampling ratio $r=n^{0.9}$.
For Aim (ii), we assess our method’s performance across different nonlinear forms of $g$ in three regression frameworks. Unlike the basic settings, we modify $g$ to examine how different nonlinear generative models impact estimation accuracy and to demonstrate ESM’s adaptability. Specifically, we define
$g_{1}(\mathbf{x})=x_{1}+|x_{2}x_{3}|-e^{0.2x_{4}}-\log(1+2|x_{5}|),$
which introduces interactions, exponential decay, and logarithmic effects. Compared to $g$, $g_{1}$ skews logistic model probabilities toward 0 and also produces a narrower range for Poisson model parameters $\lambda$.
Table 2 shows that results under $g_{1}$ remain consistent with previous findings. Increasing sample size decreases MAE for $\psi^{\prime}(\widehat{f}^{B})$ and $\widehat{f}^{B}$. Empirical standard deviations (empSD) and estimated standard errors (SE) remain well aligned, with coverage probabilities maintaining 95% validity. ESM also achieves narrower confidence intervals without sacrificing coverage accuracy compared to bootstrap intervals. Figure 2 illustrates histograms of true versus estimated transformed predictions for $g_{1}(\mathbf{x})$. Unlike Figure 1, where predictions skew toward 1, the logistic model’s $p(\mathbf{x})$ under $g_{1}$ shifts leftward toward 0. In the Poisson model, $\lambda(\mathbf{x})$ spans a narrower range than under $g(\mathbf{x})$, demonstrating ESM’s robustness across these two nonlinear settings.
Table 3: Simulation results for different models under varying sample sizes $n$ and subsampling ratios $r$ with $p=100$.
$\text{Bias}_{f}$ $\text{MAE}_{f}$ $\text{Bias}_{\psi^{\prime}}$ $\text{MAE}_{\psi^{\prime}}$ EmpSD SE CP AIL
@@
$n=400,r=n^{0.8}$0.66 0.88(0.57) -0.90 0.99(0.89) 1.05 0.72 84.5% 2.38(0.97)
-0.67 0.89(0.60) -0.85 0.96(0.86) 1.07 0.85 92.1% 3.16(1.41)
-0.64 0.87(0.63) -0.77 0.91(0.83) 1.08 0.97 95.5% 4.26(2.06)
Boostrap with $f_{0}$0.60 1.07(0.90) -0.52 1.05(0.96) 1.40 - 95.5% 21.9(22.2)
$n=700,r=n^{0.8}$0.67 0.89(0.63) -0.81 0.93(0.82) 1.09 0.75 87.2% 2.50(1.07)
$r=n^{0.85}$0.61 0.86(0.65) -0.72 0.88(0.78) 1.08 0.83 91.4% 3.19(1.43)
$r=n^{0.9}$0.52 0.82(0.65) -0.61 0.84(0.74) 1.05 0.92 94.0% 4.11(1.93)
Boostrap with $f_{0}$0.50 0.94(0.85) -0.44 0.92(0.84) 1.27 - 96.2% 19.9(18.9)
Binomial Model
$n=400,r=n^{0.8}$ 0.49 1.19(0.99) 0.07 0.17(0.16) 1.55 1.01 81.9% 0.49(0.01)
$r=n^{0.85}$ 0.55 1.27(1.06) 0.06 0.17(0.16) 1.65 1.20 86.0% 0.54(0.01)
$r=n^{0.9}$ 0.58 1.37(1.13) 0.05 0.17(0.16) 1.78 1.39 88.6% 0.59(0.00)
Boostrap with $f_{0}$ 0.31 1.39(1.08) 0.02 0.18(0.17) 1.76 - 95.9% 0.77(0.24)
$n=700,r=n^{0.8}$ 0.68 1.12(0.95) 0.09 0.15(0.15) 1.47 1.11 88.1% 0.50(0.00)
$r=n^{0.85}$ 0.69 1.16(0.98) 0.08 0.14(0.14) 1.52 1.25 90.8% 0.54(0.00)
$r=n^{0.9}$ 0.68 1.18(1.01) 0.07 0.14(0.14) 1.55 1.39 93.4% 0.59(0.01)
Boostrap with $f_{0}$ 0.51 1.33(1.05) 0.06 0.17(0.16) 1.69 - 96.5% 0.76(0.23)
For Aim (iii), we increase the number of input dimensions while retaining only a few elements that are relevant to the true output logits $f_{0}$. As the dimensionality grows, the data signals become increasingly obscured within the larger feature space, making it harder for neural networks to accurately capture these signal features. This results in a significant deviation between the estimated $\mathbb{E}(y|\mathbf{x})$ and the true $\mathbb{E}(y|\mathbf{x})$, which can also lead to inflated overall interval lengths. Table 3 shows that $\text{MAE}_{f}$, $\text{MAE}_{\psi^{\prime}}$, and AIL are notably larger compared to those in Table 1. Despite these challenges, the inference framework remains remarkably robust. The coverage probability (CP) stays stable at around 95% in most cases, even under large deviations, except when the subsample size $r$ is relatively small.
Table 4: Simulation results for different models under varying sample sizes $n$ and subsampling ratios $r$ with a deeper neural network.
$\text{Bias}_{f}$ $\text{MAE}_{f}$ $\text{Bias}_{\psi^{\prime}}$ $\text{MAE}_{\psi^{\prime}}$ EmpSD SE CP AIL

### v3

@@ A.4 Additional Experiments on Dimensionality and Nonlinearity
$n=400,r=n^{0.8}$ -0.60(0.49) 0.63(0.45) -0.30(0.29) 0.33(0.26) 0.43 0.53 0.44 77.1% 0.90(0.54)
$r=n^{0.85}$ -0.49(0.49) 0.55(0.42) -0.25(0.31) 0.30(0.26) 0.44 0.58 0.48 86.6% 1.11(0.66)
$r=n^{0.9}$ -0.37(0.51) 0.49(0.40) -0.17(0.35) 0.29(0.26) 0.48 0.62 0.51 90.7% 1.37(0.85)
$n=700,r=n^{0.8}$ -0.46(0.39) 0.49(0.35) -0.25(0.26) 0.28(0.23) 0.35 0.48 0.35 78.3% 0.79(0.47)
$r=n^{0.85}$ -0.31(0.38) 0.38(0.30) -0.18(0.27) 0.24(0.22) 0.35 0.50 0.37 88.5% 0.95(0.56)
$r=n^{0.9}$ -0.20(0.38) 0.33(0.27) -0.11(0.30) 0.23(0.22) 0.36 0.53 0.38 92.8% 1.11(0.68)
$n=2000,r=n^{0.8}$ -0.22(0.26) 0.27(0.20) -0.14(0.20) 0.19(0.17) 0.24 0.46 0.25 86.6% 0.67(0.39)
$r=n^{0.85}$ -0.15(0.27) 0.24(0.19) -0.10(0.22) 0.17(0.17) 0.26 0.49 0.26 91.7% 0.78(0.46)
$r=n^{0.9}$ -0.08(0.26) 0.22(0.17) -0.05(0.24) 0.17(0.17) 0.25 0.51 0.27 94.3% 0.85(0.52)
Table S.8: Simulation results with neural networks under $p=7$.
$\text{Bias}_{f}$ $\text{MAE}_{f}$ $\text{Bias}_{\psi^{\prime}}$ $\text{MAE}_{\psi^{\prime}}$ EmpSD SE $\text{SE}_{c}$ CP AIL
@@
$n=400,r=n^{0.8}$ -0.20(0.30) 0.28(0.23) -0.12(0.22) 0.18(0.17) 0.27 0.34 0.28 90.6% 0.74(0.42)
$r=n^{0.85}$ -0.12(0.28) 0.23(0.20) -0.07(0.22) 0.16(0.16) 0.26 0.33 0.27 93.7% 0.78(0.47)
$r=n^{0.9}$ -0.07(0.26) 0.21(0.17) -0.04(0.22) 0.15(0.16) 0.25 0.31 0.26 94.6% 0.78(0.48)
$n=700,r=n^{0.8}$ -0.09(0.20) 0.17(0.15) -0.06(0.16) 0.12(0.12) 0.19 0.26 0.19 92.8% 0.55(0.32)
$r=n^{0.85}$ -0.03(0.19) 0.15(0.13) -0.03(0.16) 0.11(0.11) 0.17 0.26 0.18 93.2% 0.52(0.26)
$r=n^{0.9}$ -0.02(0.20) 0.15(0.12) -0.02(0.17) 0.12(0.12) 0.18 0.26 0.19 94.4% 0.59(0.34)
Table S.9: Simulation results with neural networks under different true function $g(\mathbf{x})$.
$\text{Bias}_{f}$ $\text{MAE}_{f}$ $\text{Bias}_{\psi^{\prime}}$ $\text{MAE}_{\psi^{\prime}}$ EmpSD SE $\text{SE}_{c}$ CP AIL